OpenStudy (anonymous):

Estimate the square root of -11 tot he nearest integer A. -4 B. -5 C. -2 D. -3

OpenStudy (mathmale):

If you truly mean\[\sqrt{-11},\] then none of the four possible answers is correct. Please double-check: Have you copied down the problem exactly as it was given?

OpenStudy (anonymous):

i think its -3 because if you put a negative in a square root its not possible

OpenStudy (anonymous):

and thats the closest to it without the negative

OpenStudy (anonymous):

\[-\sqrt{11}\]

OpenStudy (anonymous):

\[\sqrt{11}=3.3166\approx3\]

OpenStudy (mathmale):

Hint: \[\sqrt{-1}=i\]

OpenStudy (anonymous):

\[-\sqrt{11}=-3.3166\approx-3\]

OpenStudy (anonymous):

Thank you guys!

OpenStudy (mathmale):

You must take that identity into account in solving this problem. But again, I assert that none of the four possible answers are correct. You cannot simply take the negative sign out from under the radical and put it in front of Sqrt(11). Again, please note, Sqrt(-1) = i where i is the imaginary operator.

OpenStudy (mathmale):

Please go back and compare your presentation of this problem to the original problem statement. Please note that -3 is not the correct answer, because of that negative sign under the radical.

OpenStudy (anonymous):

@mathmale , @CoolKatt16 wrote it wrong the first time.. she wrote it again correctly.

OpenStudy (anonymous):

\[-\sqrt{11}\]

OpenStudy (mathmale):

@coolkatt16: This all goes to show how important it is to double check your problem presentation before you hit that Post button. Hello1213 is correct in stating that the INTERMEDIATE (not final) answer is \[-\sqrt{11}\]But you must now evaluate Sqrt(11), round that off to the nearest integer, and then present your final answer. What is it?